OpenStudy (anonymous):

Solve the equation (eqn coming soon) . . . . ANSWER = {4/13,9/13}

OpenStudy (anonymous):

drum roll....

OpenStudy (anonymous):

\[\sqrt{\frac{ 1-x }{ x }}+\sqrt{\frac{ x }{ 1-x }}=\frac{ 13 }{ 6 }\]

OpenStudy (anonymous):

anyone got it

hartnn (hartnn):

long method...

OpenStudy (anonymous):

maybe

hartnn (hartnn):

squaring twice.....

OpenStudy (anonymous):

common goku release your spirit bomb on this question :D

hartnn (hartnn):

ok, lets go for long method :P square both sides and tell me what u get ?

OpenStudy (anonymous):

wait wait wait

hartnn (hartnn):

oh, wait. its simple. u need to square only once!

hartnn (hartnn):

square one and tell me what u get......

hartnn (hartnn):

*once

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[u^2 = \frac{1-x}{x}\]

OpenStudy (anonymous):

what happened?

hartnn (hartnn):

lol. square the equation

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can we do this y + 1/y = 13/6

OpenStudy (anonymous):

yeah, that works

OpenStudy (anonymous):

do what hartnn suggests

OpenStudy (anonymous):

I believe that it is faster

OpenStudy (anonymous):

what is faster?

OpenStudy (anonymous):

I think the squaring is faster..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if we square the left side it will become a formula (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

by squaring one side it will remove the roots isn't it @hartnn

hartnn (hartnn):

yup

OpenStudy (anonymous):

is'nt it a tuff question

hartnn (hartnn):

no

OpenStudy (anonymous):

ok

hartnn (hartnn):

just algebra

jhonyy9 (jhonyy9):

@hartnn ,@Miyuru sorry just a little help so when there are sqaurroots and fractions with undefined terms like x or/and y in denominator or for radical is indifferent so than what will be the first step --- condition,restriction what we need to doing for can sure existing this radical and fractions so for not being these all undefined please make it before all !!! RIGHT this in this way please ???

jhonyy9 (jhonyy9):

so Koli this is the first step ALLWAYS that make the condition for denominator and in this case forradical that for what value of x not can being equal zero NEVER OK ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than in this case x not can being equal 0 and ???

OpenStudy (anonymous):

can we suppose the whole root thing in one variable like y

OpenStudy (anonymous):

this will make it easy

OpenStudy (anonymous):

and the other root will be like 1/y or something

OpenStudy (anonymous):

this bit difficult

OpenStudy (anonymous):

the question?

jhonyy9 (jhonyy9):

ok i understand it but i have learned in my schoole that hwen there is an exercise with radical and fractions ALLWAYS the first step is necessary being to make the existing conditions for radical and fraction so for these never not being undefined ok ?

OpenStudy (anonymous):

yes, but the radical is defined at both of the roots which are given as answers

OpenStudy (anonymous):

For me we have taught to square to remove the roots

OpenStudy (anonymous):

me too

jhonyy9 (jhonyy9):

@Miyuru ok i understand your message but indifferent what you like to do and how the first step i think that need being these because this exercise was wrote hence and just when these conditions are satisfied so just than this radical and fractions will be defined and just in this case you can solve this equation sorry but i think in this way ,this is my opinion !!!

OpenStudy (anonymous):

plz anybody write the steps on how to do this question i am getting confused

hartnn (hartnn):

(1-x)/x +x/(1-x) +2 = 169/36 did u get this after squaring?

OpenStudy (anonymous):

yes got this

hartnn (hartnn):

multiply 36x(1-x) on both sides

OpenStudy (anonymous):

done

hartnn (hartnn):

simplify to get a quadratic in x

OpenStudy (anonymous):

\[\frac{ 1-x }{ x }+\frac{ x }{ 1-x }+2*\frac{ 36x }{ 1-x }=\frac{ 169 }{ 1-x }\]

OpenStudy (anonymous):

stuck here

hartnn (hartnn):

?? u did not multiply 36x(1-x) to each term!

OpenStudy (anonymous):

\[\frac{ 1-x }{ x }+\frac{ x }{ 1-x}+2*36x(1-x)=\frac{ 169x }{ 1-x }\]

hartnn (hartnn):

what is 36x(1-x) * (1-x) / x ??

OpenStudy (anonymous):

u said multiply 36x(1-x) to both sides

OpenStudy (anonymous):

woah!!!!!!!!!

OpenStudy (anonymous):

at last your typing is finished and got the whole question solved

OpenStudy (anonymous):

@shubham.bagrecha your algebra is impressive

jhonyy9 (jhonyy9):

yes koli this is true but we here in openstudy wann learning people how need solving exercises,problems and not just adding the right answer OK ? dont forget it please @shubham.bagrecha ATTENTION !!! please

OpenStudy (anonymous):

sorry @jhonyy9 but i was confused in this type,u know i try to do all questions by myself

jhonyy9 (jhonyy9):

ok i understand it but these words i have wrote firstly to @shubham.bagrecha

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

jhonny i know that.

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

satisfied?

jhonyy9 (jhonyy9):

koli - so hope that you have understood it right all ,correct and just a little point i like to courage you that your first step with (1-x)/x =y was right sure - my opinion - so if you continue it on this way you have got these results too easy exactly the same like @shubham.bagrecha just COURAGE Koli ,OK. ???